OpenStudy (anonymous):

Factor: x2 - 9x + 12 (x + 1)(x - 12) (x - 2)(x + 6) (x - 2)(x - 6)

OpenStudy (anonymous):

last option is prime! Forgot to type it though!

OpenStudy (anonymous):

will give a medal for best response!

OpenStudy (zehanz):

The last number (12) of your trinomial should be the product of the numbers between the brackets (those on the right) First one: 1*(-12)=-12 wrong Second one: (-2)*6=-12 wrong Third one: (-2)*(-6)=12 right! Could be the third one, but there is an extra condition: the sum of the numbers should be equal to -9 (the coefficient of x) -2+-6=-8 wrong So it could only be the fourth, unknown ;) option...

OpenStudy (anonymous):

thanks!

OpenStudy (zehanz):

YW!